OpenStudy (anonymous):

"What is the solution of the following system?" -3x - 2y = -12 9x + 6y = -9 @jigglypuff314

jigglypuff314 (jigglypuff314):

for this, you can divide through by 3 for the second equation :)

OpenStudy (anonymous):

Divide the second equation by 3???

OpenStudy (anonymous):

Woot! Okay.

OpenStudy (anonymous):

I got 3x+2y=-3

jigglypuff314 (jigglypuff314):

Great! :D wow, you're getting good at this! :) now you can do this by elimination -3x - 2y = -12 3x + 2y = -3

OpenStudy (anonymous):

do i add or subtract? And I can cross out -2y and +2y right?

jigglypuff314 (jigglypuff314):

oh wait O_O when you see that by doing that you get no variables left and unequal numbers on the right, it means that the two equations are parallel

OpenStudy (anonymous):

OH ! so it's has infinitely many solutions ?

jigglypuff314 (jigglypuff314):

hmmm parallel lines would never intersect so no solutions but if the same thing happened on the left, but the right were Equal. then since they are the same equations, there would be infinitely many solutions

OpenStudy (anonymous):

Oh okay, so the left is no solutions, the right is infinitely many?

jigglypuff314 (jigglypuff314):

mmmm sorry if I was confusing u I meant that for this you were left with not variables on the left, but since -12 and -3 are different the system would be of parallel lines, which never intersect so have no solutions if you had numbers like 3 and 3 on the right then the system would be of same lines, which overlap entirely and have infinite solutions

OpenStudy (anonymous):

Okay! Gotcha, I understand! Thank you so much for all your help. If I need anything, I'll let you know. :)

jigglypuff314 (jigglypuff314):

Okay :) I'm glad I could help <3